# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v1

@@ 2.2 The proposed ML framework
The proposed methodology consists of four main steps. First, we map the discrete distributions of pedestrian positions into continuous density fields defined on a spatial grid using Kernel Density Estimation (KDE) [77]. Second, we construct a restriction operator to project the resulting density fields into an appropriate latent space. For this purpose, we employ the POD [78] method on normalized density profiles across the computational domain, ensuring that mass conservation is preserved in the reconstructed fields. Third, within this latent space, we train surrogate autoregressive ROMs using MVAR and LSTM time-series forecasting models to capture the complex dynamics present in the sequential data. Finally, in the fourth step, we lift the latent ROM dynamics back to the macroscopic density ambient space by solving the preimage problem. In the case of POD, this lifting step is computationally inexpensive, as it reduces to a simple linear projection onto the ambient space using the stored orthogonal basis. A schematic overview of the methodology is shown in Fig. 1.
Figure 1: Schematic of the proposed methodology. First, microscopic distributions of pedestrian positions are mapped to macroscopic density fields (Step 1) and embedded into a latent space using POD (Step 2). Second, surrogate reduced-order models (MVAR and LSTM) are trained in this latent space to capture and forecast the complex dynamics (Step 3). Finally, the learned dynamics are lifted back to the high-dimensional density fields via linear projection with the constructed POD basis (Step 4).
@@ 2.2.1 From discrete pedestrian positions to continuous density fields
The initial step in our framework involves constructing a continuous density field from the discrete positions of pedestrians. Given microscopic observations, obtained by agent-based simulators in the form of Eq. (1, 2) or experimental data, macroscopic continuous quantities are typically derived via local averaging at each point of the spatial domain where the crowd moves. Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator in Appendix A. We next employ Multivariate Kernel Density Estimation (KDE) [79] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
@@
where $K_{h}$ is a smooth non-negative kernel function satisfying $\int_{\Omega}K(\mathbf{x})\,d\mathbf{x}=1$. Using a standard uniform spatial discretization $\delta x=(b-a)/n_{x}$, $\delta y=(c-d)/n_{y}$ with $n_{x}$$n_{y}$ grid points along the xx/$y$-axis of the domain $\Omega$, the resulting from Eq. (6) density fields are computed as $\rho(\mathbf{x},t)=[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $\rho((x_{i},y_{j}),t)$ corresponds to the density computed at the grid point $(x_{i},y_{j})$, for $i=1,\dots,n_{x}$, $j=1,\dots,n_{y}$. Then, the total mass in the domain is given by $S(t)=\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y$. While KDE can be used for extracting continuous macroscopic fields, such as momentum and energy, from any microscopic discrete quantities, here we only construct the density fields in Eq. (6). Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid-points of the obstacle), and the treatment of periodic boundary conditions are provided in Appendix C.
@@ 2.2.2 The Restriction and Lifting Operators
Having obtained the macroscopic density fields, we now proceed to the second and fourth key elements of the approach, which, in analogy to the Equation-free framework, constitute the construction of an “embedding/restriction” and “lifting” operators to and from an appropriate low-dimensional latent space.
The restriction operator, say, $\mathcal{R}$, [35], is a map from density distributions $\rho(\mathbf{x},t)$ to a low-dimensional latent space coordinates, say $Y_{d}(t)$, i.e.,:
@@
where $n_{c}=n_{x}\times n_{y}$. To construct the restriction operator, we apply Proper Orthogonal Decomposition (POD) [78] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq. (6), for different initial conditions; see Appendix B for details. Denoting the densities by $\rho^{(c)}(\mathbf{x},t_{0}+k\delta t)\equiv\rho^{(c)}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$, where $\delta t$ is a microscopic inner simulation step/observation sampling time of an experiment/real data, $t_{0}$ is the initial time, and the superscript $(c)$ denotes the case $c=1,\ldots,C$ of the different initial conditions considered, the training dataset generated by microscopic runs is represented by the matrix $X\in\mathbb{R}^{n_{c}\times(C\cdot K)}$:
@@
with columns, say, $x_{k}\in\mathbb{R}^{n_{c}}$ being the normalized density field along cases and timesteps $k=1,\ldots,n_{t}=C\cdot K$.
@@
$\bar{X}=XH=U\Sigma V^{T},\quad H=I_{n_{t}}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top},$ (10)
where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular vectors, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim. unity matrix and $\mathbf{1}_{n_{t}}$ is the $n_{t}$-dim. unit column vector. The latent space is spanned by the first $d$ left-singular vectors forming $U_{d}$. Thus, projecting $X$ on the latent space we get, the low-dimensional embedding:
$Y_{d}=U_{d}^{T}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}$ (11)
For the lifting operator that maps $Y_{d}$ to the space of density profiles, we demonstrate that the POD reconstruction (aka the POD lifting operator) preserves the original density. This aligns with the general characteristic of POD in maintaining symmetries [80]. Here, in particular, we demonstrate the following proposition.
@@
Let us assume a matrix $X\in\mathbb{R}^{n_{c}\times n_{t}}$, with columns being normalized density fields. If the density is preserved along timesteps, i.e., if
@@
then the reconstructed density field computed by the POD as:
@@
where $U_{d}=[u_{1},u_{2},\dots u_{d}]\in\mathbb{R}^{n_{c}\times d}$ is the orthonormal basis formed by the first $d$ left-singular vectors computed by the SVD, is also preserved, i.e.,:
@@
As by hypothesis, the sum of each column of $X$ is equal to one (that is $\mathbf{1}_{n_{c}}^{\top}X=\mathbf{1}_{n_{t}}^{\top}$), we have that:
@@
This implies that $\mathbf{1}_{n_{c}}$ is orthogonal to the column space of $\bar{X}$, and therefore is orthogonal to the left singular vectors of the SVD decomposition of $\bar{X}$, which provide an orthogonal base for the column space, i.e.,:
@@
where $d=1,2,\dots r$, with $r=rank(X)$. Then, the multiplication of the reconstructed matrix in Eq. (13) by $\mathbf{1}_{n_{c}}^{\top}$ implies:
@@
$\mathbf{1}_{n_{c}}^{\top}\tilde{X}=\begin{bmatrix}\sum_{i=1}^{n_{c}}\hat{x}_{i,1}&\sum_{i=1}^{n_{c}}\hat{x}_{i,2}&\dots&\sum_{i=1}^{n_{c}}\hat{x}_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}}.$ (18)
Hence, the POD reconstruction preserves the original density. ∎
@@
$x(t)=\dfrac{1}{\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y}[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{c}}$ (19)
and define the restriction operator $\mathcal{R}$ by the projection on the POD basis of Eq. (11) as:
@@ 2.2.3 Learning the dynamics in latent spaces using MVAR and LSTM
With the restriction operator constructed, which maps density fields to latent coordinates, the third step of the proposed approach involves training ROMs to learn the latent dynamics from sequential data. We note that while the restriction and lifting operators $\mathcal{R}$ and $\mathcal{L}$ are, in principle, defined over continuous time inputs, we apply them at a discrete set of uniformly spaced time instances $t_{k}=t_{0}+k\delta t$ for $k=1,2,\ldots,K$, consistent with the temporal sampling of our data; see Eq. (8). To consider lags in the latent coordinate embeddings $\boldsymbol{y}(t_{k})$, we employ both linear and nonlinear autoregressive models, namely MVARs and LSTMs.
@@
$\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})]=\mathbf{0},\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{k})^{\top}]=\Sigma,\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{r})^{\top}]=\mathbf{0},\quad\text{for }t_{k}\neq t_{r},$ (23)
@@
for $c=1,\ldots,C$ and $k=1,\ldots,K$. The optimization problem in Eq. (24) can be solved using methods such as regularized least square methods; in particular, we used ridge regularization [81]. We note that MVARs require the trivial assumption of time series stationarity, which was the case for our illustrations. If the latent coordinate time series is non-stationary, pre-processing steps like differencing can be applied to achieve stationarity before solving the optimization problem. We highlight that MVAR model training further requires the selection of an optimal lag window size $w$, which determines the length of the time series history used to model the temporal dependencies. Here, we used two information-theoretic criteria for determining $w$, the details for which are provided in Appendix D.
@@
where $\mathbf{h}_{t_{k}}\in\mathbb{R}^{N_{h}}$ and $\mathbf{c}_{t_{k}}\in\mathbb{R}^{N_{h}}$ are the, evaluated at discrete times $t_{k}$, hidden and memory cell states, respectively, $N_{h}$ corresponds to the number of hidden units and $\mathbf{W}_{y}\in\mathbb{R}^{d\times N_{h}},\mathbf{b}_{y}\in\mathbb{R}^{d}$ define an output layer mapping the last hidden state $\mathbf{h}_{t_{k-1}}$ to the predicted latent variable vector $\hat{\boldsymbol{y}}_{c}(t_{k})\in\mathbb{R}^{d}$. The gating mechanism of the LSTM is included in Eq. (27) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for the conciseness of the presentation; a vanilla LSTM including 4 gates, contains the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
Similarly to the sequence-to-one learning of MVARs, given sequential data for $\boldsymbol{y}_{c}(t_{k})$ for a fixed lag window of width $w$, we train the LSTM model by solving the optimization problem:
@@ 3.1 Configuration
Figure 2: The case study: The SFM simulating pedestrians moving in a corridor with an obstacle.
@@
$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{i,j}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{i,B}(\bm{x}_{i})}_{\text{repulsive force}}+\mathbf{F}_{rand},$ (31)
@@
We simulate a crowd of $N=100$ pedestrians moving through a corridor, with initial spatial distributions ranging from homogeneous dispersions to clustered configurations (for more details, see in the Appendix B). Pedestrians are directed to move from left to right by setting their initial target at coordinates $(x,y)=(25\,\text{m},\,9.6\,\text{m})$, corresponding to the midpoint of the corridor in the horizontal direction and 80% of its vertical extent. Upon reaching this intermediate waypoint, their final destination is updated to $(x,y)=(48\,\text{m},\,4.2\,\text{m})$, which lies at the right end of the corridor and at 35% of its height. To maintain a constant population density, we implement periodic boundary conditions by re-spawning agents that reach the right boundary at $x=48\,\text{m}$, at the left boundary at $x=0\,\text{m}$.
Pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator. Such configuration ensures a controlled pedestrian flow through the simulated environment and also ensures the pedestrians avoid the obstacle each time they reenter the corridor.

### v2

@@ 2.2 The proposed ML framework
The proposed methodology consists of four main steps. First, we map the discrete distributions of pedestrian positions into continuous density fields defined on a spatial grid using KDE [82]. Second, we construct a restriction operator to project the resulting density fields into an appropriate latent space. For this purpose, we employ the POD [83] method on normalized density profiles across the computational domain, ensuring that mass conservation is preserved in the reconstructed fields. Third, within the POD latent space, we train surrogate autoregressive ROMs using MVAR and LSTM time-series forecasting models to capture the complex dynamics present in the data. Finally, in the fourth step, we lift the latent ROM dynamics back to the macroscopic density ambient space by solving the pre-image problem. In the case of POD, this lifting step is computationally inexpensive, as it reduces to a simple linear projection onto the ambient space using the stored orthogonal basis. A schematic overview of the methodology is shown in Fig. 1.
Figure 1: Schematic of the proposed methodology. First, microscopic distributions of pedestrian positions are mapped to macroscopic density fields (Step 1) and embedded into a latent space using POD (Step 2). Second, surrogate reduced-order models (MVARs and LSTMs) are trained in this latent space to capture and forecast the complex dynamics (Step 3). Finally, the learned dynamics are lifted back to the high-dimensional density fields via linear projection with the constructed POD basis (Step 4).
@@ 2.2.1 From discrete pedestrian positions to continuous density fields
The initial step in our framework involves constructing a continuous density field from the discrete positions of pedestrians. Given microscopic observations, obtained by agent-based simulators in the form of Eq. (1, 2) or experimental data, macroscopic continuous quantities are typically derived via local averaging at each point of the spatial domain where the crowd moves. Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator, detailed in Appendix A. We next employ KDE [84] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
@@
where $K_{h}$ is a smooth non-negative kernel function satisfying $\int_{\Omega}K(\mathbf{x})\,d\mathbf{x}=1$. Using a standard uniform spatial discretization $\delta x=(b-a)/n_{x}$, $\delta y=(c-d)/n_{y}$ with $n_{x}$$n_{y}$ grid points along the xx/$y$-axis of the domain $\Omega$, the resulting from Eq. (6) density fields are computed as $\rho(\mathbf{x},t)=[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $\rho((x_{i},y_{j}),t)$ corresponds to the density computed at the grid point $(x_{i},y_{j})$, for $i=1,\dots,n_{x}$, $j=1,\dots,n_{y}$. Then, the total mass in the domain is given by $S(t)=\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y$. While KDE can be used for extracting continuous macroscopic fields, such as momentum and energy, from any microscopic discrete quantities, here we only construct the density fields in Eq. (6). Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid points of the obstacle), and the treatment of periodic boundary conditions are provided in Appendix C.
@@ 2.2.2 The Restriction and Lifting Operators
Having obtained the macroscopic density fields, we now proceed to the second and fourth key elements of the approach, which, in analogy to the EF framework, constitute the construction of an “embedding/restriction” and “lifting” operators to and from an appropriate low-dimensional latent space.
The restriction operator, say, $\mathcal{R}$, [39], is a map from density distributions $\rho(\mathbf{x},t)$ to low-dimensional latent space coordinates, say $Y_{d}(t)$, i.e.,:
@@
where $n_{c}=n_{x}\times n_{y}$. To construct the restriction operator, we apply POD [83] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq. (6), for different initial conditions; see Appendix B for details. Denoting the densities by $\rho^{(c)}(\mathbf{x},t_{0}+k\delta t)\equiv\rho^{(c)}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$, where $\delta t$ is a microscopic inner simulation step/observation sampling time of an experiment/real data, $t_{0}$ is the initial time, and the superscript $(c)$ denotes the case $c=1,\ldots,C$ of the different initial conditions considered, the training dataset generated by microscopic runs is represented by the matrix $X\in\mathbb{R}^{n_{c}\times(C\cdot K)}$:
@@
with columns, say $x_{k}=\rho^{(c)}(\mathbf{x},t_{k})/S(t_{k})\in\mathbb{R}^{n_{c}}$, being the normalized density fields along cases and time steps $k=1,\ldots,n_{t}=C\cdot K$.
@@
$\bar{X}=XH=U\Sigma V^{\top},\quad H=I_{n_{t}}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top},$ (10)
where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular vectors, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim. unity matrix and $\mathbf{1}_{n_{t}}$ is the $n_{t}$-dim. unit column vector. The latent space is spanned by the first $d$ left-singular vectors forming $U_{d}$. Thus, the projection of $X$ on the latent space provides the low-dimensional embedding:
$Y_{d}=U_{d}^{\top}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}$ (11)
For the lifting operator that maps $Y_{d}$ to the space of density profiles, we demonstrate that the POD reconstruction (aka the POD lifting operator) preserves the original total density. This aligns with the general characteristic of POD in maintaining symmetries [85]. Here, in particular, we demonstrate the following proposition.
@@
Let us assume a matrix $X\in\mathbb{R}^{n_{c}\times n_{t}}$, with columns being normalized density fields. If the density is preserved along time steps, i.e., if:
@@
then, the reconstructed density field, computed by the POD as:
@@
where $U_{d}=[u_{1},u_{2},\dots u_{d}]\in\mathbb{R}^{n_{c}\times d}$, is the orthonormal basis formed by the first $d$ left-singular vectors computed by the SVD, is also preserved, i.e.,:
@@
By construction, since the sum of each column of $X$ is equal to one (that is $\mathbf{1}_{n_{c}}^{\top}X=\mathbf{1}_{n_{t}}^{\top}$), we have that:
@@
This implies that $\mathbf{1}_{n_{c}}$ is orthogonal to the column space of $\bar{X}$, and therefore is orthogonal to the left singular vectors of $\bar{X}$, which provide an orthogonal basis for the column space, i.e.,:
@@
where $d=1,2,\dots r$ with $r=rank(X)$. Then, the multiplication of the reconstructed matrix in Eq. (13) by $\mathbf{1}_{n_{c}}^{\top}$ implies:
@@
$\mathbf{1}_{n_{c}}^{\top}\tilde{X}=\begin{bmatrix}\sum_{i=1}^{n_{c}}\tilde{x}_{i,1}&\sum_{i=1}^{n_{c}}\tilde{x}_{i,2}&\dots&\sum_{i=1}^{n_{c}}\tilde{x}_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}}.$ (18)
Hence, the POD reconstruction preserves the total original density. ∎
@@
$x(t)=\dfrac{1}{S(t)}[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{c}},$ (19)
where $S(t)$ is the total mass on time $t$, and define the restriction operator $\mathcal{R}$ by the projection on the POD basis of Eq. (11) as:
@@ 2.2.3 Learning the dynamics in latent spaces using MVAR and LSTM
With the restriction operator, which maps density fields to latent coordinates, constructed, the third step of the proposed approach involves training ROMs to learn the latent dynamics from sequential data. We note that while the restriction and lifting operators $\mathcal{R}$ and $\mathcal{L}$ are, in principle, defined over continuous time inputs, we apply them at a discrete set of uniformly spaced time instances $t_{k}=t_{0}+k\delta t$ for $k=1,2,\ldots,K$, consistent with the temporal sampling of our data; see Eq. (8). To consider lags in the latent coordinate embeddings $\boldsymbol{y}(t_{k})$, we employ both linear and nonlinear autoregressive models, namely MVARs and LSTMs.
@@
$\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})]=\mathbf{0},\qquad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{k})^{\top}]=\Sigma\quad\text{and}\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{r})^{\top}]=\mathbf{0}\quad\forall k\neq r,$ (23)
@@
for $c=1,\ldots,C$ and $k=1,\ldots,K$. The optimization problem in Eq. (24) can be solved using methods such as regularized least squares; in particular, we used ridge regularization [86]. We note that MVARs require the trivial assumption of time series stationarity, which was the case for our illustrations. If the latent coordinate time series is non-stationary, pre-processing steps like differencing can be applied to achieve stationarity before solving the optimization problem. We highlight that MVAR model training further requires the selection of an optimal lag window size $w$, which determines the length of the time series history used to model the temporal dependencies. Here, we used two information-theoretic criteria for determining $w$, the details for which are provided in Appendix D.
@@
where $\mathbf{h}_{t_{k}}\in\mathbb{R}^{N_{h}}$ and $\mathbf{c}_{t_{k}}\in\mathbb{R}^{N_{h}}$ are the evaluated at discrete times $t_{k}$, hidden and memory cell states, respectively, $N_{h}$ corresponds to the number of hidden units and $\mathbf{W}_{y}\in\mathbb{R}^{d\times N_{h}},\mathbf{b}_{y}\in\mathbb{R}^{d}$ define an output layer mapping the last hidden state $\mathbf{h}_{t_{k-1}}$ to the predicted latent variable vector $\hat{\boldsymbol{y}}_{c}(t_{k})\in\mathbb{R}^{d}$. The gating mechanism of the LSTM is included in Eq. (27) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for conciseness of the presentation; a vanilla LSTM including 4 gates, and contains the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
Similarly to the sequence-to-one learning of MVARs, given sequential data for $\{\boldsymbol{y}_{c}(t_{k})\}$ for a fixed lag window of width $w$, we train the LSTM model by solving the optimization problem:
@@ 3.1 Configuration
Figure 2: The case study: The SFM simulating pedestrians moving in a corridor from the left to the right with an obstacle.
@@
$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{ij}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{iB}(\bm{x}_{i})}_{\text{repulsive force}}+\mathbf{F}_{rand},$ (31)
@@
We simulate a crowd of $N=100$ pedestrians moving through a corridor, with initial spatial distributions ranging from homogeneous dispersions to clustered configurations (for more details, see in the Appendix B). Pedestrians are directed to move from left to right by setting their initial target at coordinates $(x,y)=(25\,\text{m},\,9.6\,\text{m})$, corresponding to the midpoint of the corridor in the horizontal direction and 80% of its vertical extent. Upon reaching this intermediate waypoint, their final destination is updated to $(x,y)=(48\,\text{m},\,4.2\,\text{m})$, which lies at the right end of the corridor and at 35% of its height. To maintain a constant population density, we implement periodic boundary conditions by re-spawning agents at the left boundary at $x=0\,\text{m}$, when agents reach the right boundary at $x=48\,\text{m}$.
Pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator. Such a configuration ensures a controlled pedestrian flow through the simulated environment and also ensures the pedestrians avoid the obstacle each time they reenter the corridor.